Otis Taylor has a box of coins that he uses when playing poker with his friends. The box currently contains 44 coins, consisting of pennies, dimes, and quarters. The number of pennies is equal to the number of dimes and the total value is $4.37. How many of each denomination of coin does he have in the box.
the anwser is 17 pennies, 17 dimes and 10 quarters but i need to be shown how to get those anwsers
Thanks

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

Otis Taylor has a box of coins that he uses when playing poker with his friends.
The box currently contains 44 coins, consisting of pennies, dimes, and quarters.
The number of pennies is equal to the number of dimes and the total value is $4.37.
How many of each denomination of coin does he have in the box.
--------------------
Let # of pennies be "x" ; value is x cents
# of dimes is "x" ; value is 10x cents
# of quarters is "44-2x" ; value is 25(44-2x)= 1100-50x cents
-------------------------
EQUATION:
value + value + value = 437 cents
x + 10x + 1100-50x = 437
1100 -39x = 437
39x = 663
x = 17 (# of pennies; also # of dimes)
44-2x= 44-34 = 10 (# of quarters)
===================
Cheers,
Stan H.

Answer by checkley71(8403)

P+D+Q=44
P=D THUS WE HAVE 2D+Q=44 OR Q=44-2D
.01D+.10D+.25Q=4.37 NOW REPLACE (44-2D) FOR Q
.01D+.10D+.25(44-2D)=4.37
.11D+11-.5D=4.37
-.39D=4.37-11
-.39D=6.63
D=-6.63/-.39
D=17 ANSWER FOR THE NUMBER OF PENNIES & NUMBER OF DIMES.
17=17+Q=44
Q=44=34
Q=10
PROOF
.01*17+.1*17+.25*10=4.37
.17+1.70+2.50=4.37
4.37=4.37